m4-13m2+36=0
hint x=m2
Im confused and need help

Answer by RAY100(1637)

m^4 -13m^2 +36 =0
.
Let x= m^2
.
x^2 -13x +36 =0
.
(x-9)(x-4) =0
.
x=9,,,,,,x=4
.
but x=m^2
.
m^2 =9,,,,,m= +/- 3
.
m^2 = 4,,,,,m= +/- 2
.
check,,(+/-3),,,3^4 -13(3^2) +36 =0,,,,,ok
.
(+/- 2),,,,2^4 -13(2^2) +36 =0,,,,,ok
.